4. Essay; show all work. Add the following polynomials:
(-1/3x^2-1/4x+3/4)+(1/2x^2-1/4x-1/2)
I do not understand how to do this !!!

Answer by rfer(16322)

combine like terms
-1/3x^2+1/2x^2=-2/6x^2+3/6x^2=1/6^2
-1/4x-1/4x=-2/4x=-1/2x
3/4-1/2=3/4-2/4=1/4
answer
1/6^2-1/2x+1/4